Operations Report – April 18th

Crew 210 Operations Report 18-APR-2019

SOL: 6

Name of person filing report: Shannon Rupert

Non-nominal systems: nothing to report

Notes on non-nominal systems: nothing to report

Generator (hours run last night): 10.5

On (last night): 9:30 pm

Off (this morning): 7 am

Solar— SOC 58 % It fell to 41% this afternoon but then rose again so rather than turn on the generator, I am watching it to see how it does the rest of the evening.

Diesel Reading – 36 %

Propane Reading – 49%

Ethanol Free Gasoline – 4 gallons.

Water (auxiliary tank) – 0 gallons

Water (static tank) – 475 gallons

Auxiliary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons.

Water (loft) – Static to Loft Pump used – yes

Water Meter: not read

Toilet tank emptied: yes

Deimos rover used: Still non-operational
Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes

Hours: Did not record

Beginning charge: 100

Ending charge: 62

Currently charging: yes

Spirit rover used: yes

Hours: did not check

Beginning charge: 100

Ending charge: did not record

Currently charging: yes

Opportunity rover used: no

Hours: not recorded

Beginning charge: 100

Ending charge: 94

Currently charging: yes

Curiosity rover used: yes

Hours: did not record

Beginning charge: 100

Ending charge: 64

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) none

Reason for use: n/a

Oil Added? No

ATV Fuel Used: 0 Gals.

# Hours the ATVs were Used today: 0

Notes on ATVs: nothing to report

HabCar used and why, where? Yes, to town for water

CrewCar used and why, where? Yes, south for live outreach event and then to town to pick things up at post office and have lunch

General notes and comments: We added one of the submersible pump to the transfer tank in the Hab Car but have yet to test it.

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: David will send a full report today.

Summary of ScienceDome operations: We are analyzing our samples in the lab. We collected 60 new plants this week.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing tonight, thank you.